Solve the given integral equation or integro-differential equation for y(t). y(t) + 125 | ce–vy(v) dv=8t, y(o)=0 y(t) =
Answer #1

t sely Given Integral eqh Y (t) + 125 S €-u) you) dx = 8+ w 46 ) = 0 - (2) Convelation of t* y(t) Taking Laplace tramfarm ba→ 2 (SYS) –46) +125() 40- 854 ( 18602 590-46 l{464)] = y(s), 1843 = SYS) +125 125 (1) YCS)] -46) 864 = (stie + 125(3)] YO - w→ Y(t) = 8 tlsts you a falar qurs) - 56 () es -st e 75 sto St Sin (15134 13 J5 (61 (SP) 75 -st te 75 5.53 9 (1) 853 Sin 75 Ca